OpenStudy (anonymous):

how do we solve: 4-(x-1)^2?

OpenStudy (anonymous):

..

OpenStudy (anonymous):

by "solve" do you mean "expand and combine like terms?" this is not an equation, so there is nothing to "solve"

OpenStudy (anonymous):

\[(xx-1)^2=(x-1)(x-1)=x^2-x-x+1=x^2-2x+1\] is a start therefore \[4-(x-1)^2=4-(x^2-2x+1)=4-x^2+2x-1\]

OpenStudy (anonymous):

combine like terms and get \[-x^2+2x+3\]